OpenStudy (anonymous):

Solve. 3/k+3 = 8/k-2 k = -1 k = 1 k = -6 k = 6

OpenStudy (anonymous):

@netlopes1

OpenStudy (michele_laino):

if we multiply both sides by this quantity: \((k+3)(k-2)\) we get: \[\frac{3}{{\left( {k + 3} \right)}} \cdot \left( {k + 3} \right) \cdot \left( {k - 2} \right) = \frac{8}{{\left( {k - 2} \right)}} \cdot \left( {k + 3} \right) \cdot \left( {k - 2} \right)\] please simplify